OpenStudy (anonymous):

reduce to slope-intercept form : 3x-4y=-24

OpenStudy (anonymous):

slope intercept form is y = mx + b where m is the slope, and b is the intercept. Simply isolate y on one side: -4y = -3x - 24 y = (-3/-4)x + (-24/-4) y = 3/4x + 6

OpenStudy (anonymous):

Make it y=mx+c !